# SARatrX: Towards Building A Foundation Model for SAR Target Recognition

### v2

@@ IV SARatrX Experiments
TABLE III: Comparisons of various target features for MIM. Many target features were found to be unsuitable for the multiplicative speckle noise in SAR images. This result inspired us to pursue different gradient features for SAR SSL. The pre-training settings in SSL-SAR and the model backbones formed the base version, while average velocity served as the classification metric. Bolded text indicates the best result, while underlined text is the next best result.
@@ IV-A Comparison of Model Backbones
Table II compares different model backbones used for SAR ATR, including ConvNeXt-V2 [80], ViT [79], and HiViT [41]. ConvNeXt-V2 and ViT represent the two primary architectures: a CNN and a transformer. Specifically, HiViT combines ViT with a Swin transformer [92]. The results indicated that ViT outperformed ConvNeXt-V2. On the one hand, ViT is more flexible than ConvNeXt-V2 for learning contextual information in SAR images. On the other hand, multiple downsampling steps in ConvNeXt-V2 resulted in the loss of small targets, while ViT maintained the same spatial resolution in different layers. HiViT, a visual transformer with ViT’s flexibility and hierarchical representation, outperformed ViT and ConvNeXt-V2 in our self-supervised experiments. In particular, HiViT employed small (4 × 4) input patches, capturing small target features well. Fig. 4 demonstrates that HiViT also offered a superior variable attention distance compared to ViT, due to the small target information common in remote sensing. Therefore, we use HiViT as backbone of our SARatrX.
@@ IV-B Strategy of two-step pre-training
Here, we discuss the two-step pre-training strategy included to make full use of available model weights and SAR datasets. Table II includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR. SL-ImageNet was pre-trained on ImageNet using supervised learning [^2] We used open-source weights from GitHub to conduct the experiments. Supervised weights in ConvNeXt-V2 and HiViT were obtained with supervised fine-tuning after SSL.; SSL-ImageNet was pre-trained on ImageNet from scratch using SSL; SSL-SAR was pre-trained from scratch on our SAR pre-training dataset; SSL-ImageNet & SAR pre-trained the model on a SAR dataset based on initialized weights from SSL-ImageNet.
Notice the additional supervised information introduced by SL-ImageNet did not necessarily improve SAR ATR performance (e.g., the linear probing performance of SL-ImageNet for HiVit was lower than that of SSL-ImageNet). SSL-SAR achieved better results than SSL-ImageNet using less data (12%), reflecting large differences in target features between the two images. However, ImageNet pre-training weights did provide a good initialization for lower features, such as shape and texture in SSL, with visible spectral remote sensing [93] and medical images [94]. Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table II) and attention diversity (see Fig. 4). As such, SARatrX employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
@@ IV-C Design of target signals for SAR images
After investigating the backbone and strategy, we focused on target features for SSL methods used with SAR images. One key point for MIM is designing high-quality guide signals due to the unique multiplicative speckle noise in SAR images. This means we need to suppress noise and enhance target features. As seen in Table III, we considered five target features (pixel values) [81], including a low pass filter [90], HOG features [43], deep features [91], SAR-HOG [85], and gradient by ratio [86, 85]). All SSL methods used the SSL-SAR setting and base version. We first considered whether existing methods based on ViT were suitable for SAR image classification. PixMIM [90] applies a low pass filter to remove high-frequency components, driving the model to focus on shape information. However, PixMIM did not outperform MAE because the noise type in SAR is multiplicative, and the filter parameters require a trade-off between the target and noise. FG-MAE [43] uses HOG to capture SSL features in SAR scene-level tasks, though we found that HOG did not ensure accurate SAR target features. Target regions typically exhibited strong scattering values, and the included speckle noise often caused the gradient computations to exhibit strong false points in these regions. In addition, I-JEPA [91] proposed deep networks for use as target feature encoders to capture deep semantic features. However, this can lead to training overfitting noise and a failure to learn effective feature representations.
As such, we chose SAR features as target features to enhance HiViT. SAR-HOG changes the gradient calculations for HOG features and uses gradient by ratio to solve for the speckle noise, thereby outperforming pixel values and HOG. Inspired by PixMIM, we prefer to directly use the target shape (i.e., gradient features) as a target feature  [^3] SAR-HOG uses the same multi-scale settings to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal.. In addition, multi-scale methods can improve the feature representations of various small targets common in remote sensing. Discussions of kernel settings used for computing gradients are illustrated in Fig. 3. Scale can also affect feature quality, as a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large targets and noise suppression. Therefore, combining features of different scales (see Fig. 3) offers improvements over a single scale for various image target sizes.
@@ IV-D Analysis
As stated above, the primary contributions of SARatrX can be summarized as follows: the HiViT architecture avoids the loss of small target information. SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capabilities. MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images. By taking advantage of these insights, SARatrX can learn high-quality target features from noisy SAR remote sensing images, as seen in Table II. In the next section, we analyze the diversity and scalability of SARatrX.
@@
Figure 4: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colours represent different layers for better visualization) in the SSL models. Attention distance represents the range of a receptive field. We focused specifically on model architectures (Fig. (a) v.s. Fig. (b)), initialization weights (Fig. (a) v.s. Fig. (c)), and SSL signals (Fig. (d) v.s. Fig. (e)) to ensure diverse attention ranges for SAR target recognition, including the HiViT architecture, ImageNet weights, and SAR target features.
@@
Figure 5: The scalability of SARatrX for dataset size, parameters, and training epochs with linear probing performance. While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting, due to the dataset size and speckle noise.
@@
Visualization - Prior research [95] has shown that supervised pre-training and contrastive learning only model global information in higher layers, while MIM can model both local and global information. However, we observed that this effect was not only related to the chosen method but also to the data properties. Fig. 4 (a) demonstrates that ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM modeling properties. As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure, as seen in Fig. 4 (b). In addition, using ImageNet weights for initialization also solved this problem, as seen in Fig. 4 (c). Similarly, Fig. 4 (d) demonstrates that HOG features enhanced the noise interference, which limited feature diversity. MGF effectively extracts shape information from targets, focusing the model on diverse edge information in the lower layer. However, this approach removes textures and preserves edge information, which motivates higher layers to rely less on texture details and diminishes the attention range shown in Fig. 4 (e). Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training, as seen in Fig. 4 (f).
Scaling experiment - Although MIM learns effectively and scales with data and model resources [96], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR. Fig. 5 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs. Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed significant rising curve in downstream task performance with increasing data and parameter quantities in Figs. 5 (a) and (b). This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals. However, as in [96], the model tended to overfit during extended training epochs when the pre-training set contained at least 100,000 images. In addition, SAR image noise and low resolution further aggravated the overfitting. Regardless, SARatrX outperformed our previous study (SAR-JEPA), which overfit at 400 epochs with 94,776 SAR images. Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
V Leveraging SARatrX for Recognition
We have discussed different aspects of SARatrX, but there are several datasets and specialized models available for SAR ATR. Therefore, we compared our proposed SARatrX algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [97] and PD [98]), semi-supervised learning (EUAPS [99]), and self-supervised learning (MSFA [21] and BIDFC [19]). We focused on SAR recognition tasks, including image classification and detection. More detailed settings [^4] We removed the test sets of downstream tasks from the pre-training set samples. are provided in Appendix B.
@@ V Leveraging SARatrX for Recognition
SARatrX 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
@@
Classification task - Table IV details the performance of SARatrX for the MSTAR [18] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs). Notice that SSL (BIDFC and ours) and semi-supervised models (EUAPS) significantly outperformed other methods for small samples with additional unlabeled data. Our results also surpassed the previous best by large margins. For example, SOC 1-shot accuracy increased by 4.5% and EOCs 1-shot accuracy increased by 15.1% on average. demonstrating the value of FMs in an era of rapidly growing SAR data. In particular, SARatrX exhibited a robustness to the EOC setting of variable imaging conditions. This result indicated that the foundation model could learn stable features and relationships from diverse imaging conditions in a large number of samples.
@@
Detection task - As illustrated in Table V, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft). SARatrX outperformed our previous MSFA by 0.8 points on SARDet-100K. MSFA also includes complex training processes and target features, which employ multi-stage training between RGB and SAR images with three different target features (HOG [83], Haar-like [84], WSTG [115]), including a detection pre-training step. SARatrX is thus simpler yet more effective for SAR images. Notably, SARatrX outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table V.
@@
Of course, our study is only a preliminary exploration of SSL for SAR image interpretation and more effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties. Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation include target recognition, scene classsification, semantic segmentation, and change detectionn, but this will require additional research.
@@ VI Conclusion and Future Perspectives
This study proposed SARatrX for SAR ATR and systematically investigated a foundation model framework. First, a pre-training dataset SARDet-180K was constructed from 14 open-source datasets, including various targets, scenes, and sensors. The foundation model’s pre-training backbone, SSL methods, and downstream tasks were then discussed in detail. Importantly, SARatrX demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field. We believe that further research on SAR foundation models, including SARatrX, has the potential to generalize feature representations of SAR images and accelerate all-day, all-weather target recognition in Earth observations. However, foundation model research requires large data, which is problematic for SAR images, as they are expensive and require specific imaging equipment and algorithms. Privacy and security may also prevent the data from becoming open source. Therefore, we are particularly grateful to the publishers of open-source SAR target datasets. By making SARatrX publicly available, we aim to accelerate the progress of FMs in SAR target recognition by enabling researchers to use our dataset and code to design better methods or explore downstream applications.
Although this work systematically investigated a foundation model framework, several limitations and challenges will require exploration in a future study. The SAR images were derived from open-source SAR datasets, and the targets primarily included vehicles, ships, aircraft, oil tanks, etc. Thus, collecting target samples from increasingly unlabeled imagery could further expand the amount of data to reach a million level and the range of downstream applications. In terms of model architecture, learning the dynamic contextual information on space-time is also an important issue. In addition to target shape features, various scenarios and target signatures for SAR self-supervised learning need to be explored. Finally, investigating expert knowledge with text used for multimodal interactions, and describing the relationship between targets and scenes, could further enhance the representation capabilities of the foundation model for SAR ATR.
In conclusion, we have demonstrated the ability of SARatrX to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks. By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to a generalized SAR target recognition.
Appendix A Implementation Details for Section 4 and SARatrX Experiments

### v3

@@ IV SARATR-X Experiments
TABLE V: Comparisons of various target features for MIM. Many target features were found to be unsuitable for the multiplicative speckle noise in SAR images. These results inspired us to pursue different gradient features for SAR SSL. The pre-training settings in SSL-SAR and the model backbones formed the base version, while average velocity served as the classification metric. Bolded text indicates the best result, while underlined text is the next best result.
@@ IV-A Comparison of Model Backbones
Table IV compares different model backbones used for SAR ATR, including ConvNeXt-V2 [86], ViT [85], and HiViT [40]. The results indicated that ViT outperformed ConvNeXt-V2. On the one hand, ViT is more flexible than ConvNeXt-V2 for learning contextual information in SAR images. On the other hand, multiple downsampling steps in ConvNeXt-V2 resulted in the loss of small targets, while ViT maintained the same spatial resolution in different layers. HiViT outperformed ViT and employed small (4 × 4) input patches, capturing small target features well. Fig. 5 demonstrates that HiViT also offered a superior variable attention distance compared to ViT due to the small target information common in remote sensing. Therefore, we use HiViT as the backbone of our SARATR-X.
@@ IV-B Strategy of two-step pre-training
Here, we discuss the two-step pre-training strategy included to make full use of available model weights and SAR datasets. Table IV includes four pre-training settings: SL-ImageNet, SSL-ImageNet, SSL-SAR, and SSL-ImageNet & SAR. SL-ImageNet was pre-trained on ImageNet using supervised learning [^3] We used open-source weights from GitHub to conduct the experiments. Supervised weights in ConvNeXt-V2 and HiViT were obtained with supervised fine-tuning after SSL.; SSL-ImageNet was pre-trained on ImageNet from scratch using SSL; SSL-SAR was pre-trained from scratch on our SAR pre-training dataset; SSL-ImageNet & SAR pre-trained the model on a SAR dataset based on initialized weights from SSL-ImageNet.
Notice the additional supervised information introduced by SL-ImageNet did not necessarily improve SAR ATR performance (e.g., the linear probing performance of SL-ImageNet for HiVit was lower than that of SSL-ImageNet). SSL-SAR achieved better results than SSL-ImageNet using less data (12%), reflecting large differences in target features between the two images. However, ImageNet pre-training weights did provide a good initialization for lower features, such as shape and texture in SSL, with visible spectral remote sensing [98] and medical images [99]. Our experiments also confirmed that using SSL-ImageNet as initialization weights improved the pre-training performance of SAR images (see Table IV) and attention diversity (see Fig. 5). As such, SARATR-X employed the SSL-ImageNet & SAR settings to complement the richness of pre-training.
@@ IV-C Design of target signals for SAR images
After investigating the model backbone and learning strategy, we focused on target features for SSL methods used with SAR images. One key point for MIM is designing high-quality guide signals due to the unique multiplicative speckle noise in SAR images. This means we need to suppress noise and enhance target features. As seen in Table V, we considered five target features (pixel values) [87], including a low pass filter [96], HOG features [43], deep features [97], SAR-HOG [91], and gradient by ratio [92, 91]). All SSL methods use the SSL-SAR setting and base version. We first considered whether existing methods based on ViT were suitable for SAR image classification. PixMIM [96] applies a low pass filter to remove high-frequency components, driving the model to focus on shape information. However, PixMIM did not outperform MAE because the noise type in SAR is multiplicative, and the filter parameters require a trade-off between the target and noise. FG-MAE [43] uses HOG to capture SSL features in SAR scene-level tasks, though we found that HOG did not ensure accurate SAR target features. Target regions typically exhibited strong scattering values, and the included speckle noise often caused the gradient computations to exhibit strong false points in these regions. In addition, I-JEPA [97] proposed deep networks for use as target feature encoders to capture deep semantic features. However, this can lead to training overfitting noise and a failure to effective features.
As such, we chose SAR features as target features to enhance HiViT. SAR-HOG changes the gradient calculations for HOG features and uses gradient by ratio to solve for the speckle noise, thereby outperforming pixel values and HOG. Inspired by PixMIM, we prefer to directly use the target shape (i.e., gradient features) as a target feature  [^4] SAR-HOG uses the same multi-scale settings to illustrate that simple gradient features can effectively represent target shape as an SSL guide signal.. In addition, multi-scale methods can improve the feature representations of various small targets common in remote sensing. Discussions of kernel settings used for computing gradients are illustrated in Fig. 4. Scale can also affect feature quality, as a smaller scale is finer for small target edge extraction, while a larger scale is more suitable for large targets and noise suppression. Therefore, combining features of different scales (see Fig. 4) offers improvements for various image target sizes.
@@ IV-D Analysis
As stated above, the primary contributions of SARATR-X can be summarized as follows: the HiViT architecture avoids the loss of small target information. SSL-ImageNet & SAR use ImageNet pre-training weights to provide a good initialization for diversity perceptual capabilities. MGF ensures high-quality target features and suppresses speckle noise under SSL with SAR images. By taking advantage of these insights, SARATR-X can learn high-quality target features from noisy SAR remote sensing images, as seen in Table IV. In the next section, we analyze the diversity and scalability of SARATR-X.
@@
Figure 5: Averaged attention distances for various attention heads (the x-axis is the attention head w.r.t layer number, and point colors represent different layers for better visualization) in the SSL models. Attention distance represents the range of a receptive field. We focused specifically on model architectures (Fig. (a) v.s. Fig. (b)), initialization weights (Fig. (a) v.s. Fig. (c)), and SSL signals (Fig. (d) v.s. Fig. (e)) to ensure diverse attention ranges for SAR target recognition, including the HiViT architecture, ImageNet weights, and SAR target features.
@@
Figure 6: Scalability of SARATR-X for dataset size, parameters, and training epochs with linear probing performance. While our method benefited from these three attributes, it is important to note that excessive training epochs often led to overfitting due to the dataset size.
@@
Visualization - Prior research [100] has shown that supervised pre-training and contrastive learning only model global information in higher layers, while MIM can model both local and global information. However, we observed that this effect was not only related to the chosen method but also to the data properties. Fig. 5 (a) demonstrates that ViT with MAE focuses on global information due to large SAR image scenes, which differs from MIM modeling properties. As such, HiViT includes various attention distances from 40 to 140 in different layers with its high spatial resolution and hierarchical structure but has decreased diversity in the bottom layer, as seen in Fig. 5 (b). In addition, using ImageNet weights for initialization solved this problem, as seen in Fig. 5 (c). Similarly, Fig. 5 (d) demonstrates that HOG features enhanced the noise interference, which limited feature diversity. MGF effectively extracts shape information from targets, focusing the model on diverse edge information in the lower layer. However, this approach removes textures and preserves edge information, which motivates higher layers to rely less on texture details and diminishes the attention range shown in Fig. 5 (e). Thus, we combined SSL-ImageNet &\&& SAR with MGF for two-step pre-training in Fig. 5 (f). However, the benefits of different modalities, especially a sufficient number of visible images, for SAR data need to be explored more. In the future, we plan to collect the satellite and airborne multimodal paired data and use various features, including filter pixels, handmade features, and learned features, to ensure the migration of the underlying feature extraction and the higher-level semantic information. Continuous learning [53, 101] is also needed to prevent catastrophic forgetting.
@@
SARATR-X 2024 58.7 86.1 (+5.7) 64.7 (+3.3)
Scaling experiment - Although MIM learns effectively and scales with data and model resources [121], a question arises as to whether our method can ensure scalability for MIM when dealing with noisy data, such as SAR. Fig. 6 presents the results of a scaling experiment from three perspectives: dataset size, parameters, and training epochs. Despite our pre-training set comprising 186,660 images, which is smaller than ImageNet-1K, we observed a significant rising curve in downstream task performance with increasing data and parameter quantities in Figs. 6 (a) and (b). This result indicated that the foundational model could fully achieve its potential in SAR images by extracting high-quality features as guiding signals. However, as in [121], the model tended to overfit during extended training epochs when the pre-training set contained about 100,000 images in ImageNet. In addition, SAR image noise and low resolution further aggravated the overfitting. Regardless, SARATR-X outperformed our previous study (SAR-JEPA), which overfitted at 400 epochs with 94,776 SAR images. Thus, there is a need to continue investigating new ways to ensure high-quality feature representations when extending SAR foundation models.
V Leveraging SARATR-X for Recognition
Figure 7: Visualization of detection on SARDet-100K. False alarms and missed detections are common in SAR images, especially when similar targets overlap and complex scenes are present. While our method effectively improves the detection effect by learning the contextual information in the image, target detection in complex scenes and low-quality images is still very difficult.
We have discussed different aspects of SARATR-X, but there are many datasets and specialized models available for SAR ATR. Therefore, we compared our proposed SARATR-X algorithm with other state-of-the-art techniques, such as supervised learning ($\rm{CS}^{n}$et [118] and PD [106]), semi-supervised learning (EUAPS [105]), and self-supervised learning (MSFA [39] and BIDFC [41]). We focused on SAR recognition tasks, including image classification and object detection in Fig. 2. More detailed settings [^5] We removed the test sets of downstream tasks from the pre-training set samples, and other method results are from their original article tables. are provided in Appendix B.
@@ V Leveraging SARATR-X for Recognition
Classification task - Table VI details the performance of SARATR-X for the MSTAR [36] dataset with standard operating conditions (SOCs) and extended operating conditions (EOCs). Notice that SSL (BIDFC and ours) and semi-supervised models (EUAPS) significantly outperformed other methods for small samples with additional unlabeled data. Our results also surpassed the previous best by large margins. For example, SOC 1-shot accuracy increased by 4.5% and EOCs 1-shot accuracy increased by 15.1% on average. Demonstrating the value of FMs in an era of rapidly growing SAR data. In particular, SARATR-X exhibited robustness to the EOC setting of variable imaging conditions. This result indicated that the foundation model could learn stable features and relationships from diverse imaging conditions in a large number of samples.
@@
Detection task - As illustrated in Table VII, we reported a box mAP for SAR target detection with a horizontal bounding box for multi-category (SARDet-100K and OGSOD), ship (SSDD), and aircraft detection (SAR-Aircraft). SARATR-X outperformed our previous MSFA by 0.8 points on SARDet-100K. MSFA also includes complex training processes and target features, which employ multi-stage training between RGB and SAR images with three different target features (HOG [89], Haar-like [90], WSTG [122]), including a detection pre-training step. SARATR-X is thus simpler yet more effective for SAR images. The detection visualization is shown in Fig. 7, and we have fewer missed detections and false alarm results. Notably, SARATR-X outperformed or offered comparable performance for multiple datasets, compared to several specifically designed detection methods shown in Table VII.
@@
Of course, our study is only a preliminary exploration of SSL for SAR image interpretation. More effective target features could be achieved in a data-knowledge dual-driven manner by further mining information on SAR imaging mechanisms and properties. Furthermore, given a larger dataset and computing power, the path of FMs will hopefully lead to generalized SAR interpretation, including target recognition, scene classification, semantic segmentation, and change detection, but this will require additional research.
@@ VI Conclusion and Future Perspectives
This study proposed a foundation model SARATR-X for SAR ATR. First, a pre-training dataset SARDet-180K was constructed from 14 open-source datasets, including various targets, scenes, and sensors. The foundation model’s pre-training backbone, SSL methods, and downstream tasks were then discussed in detail. Importantly, SARATR-X demonstrated superior performance on different target recognition datasets, demonstrating the potential of FMs in this field. We believe that further research on SAR foundation models, including SARATR-X, has the potential to generalize feature representations of SAR images and benefit all-day, all-weather target recognition in Earth observations. However, FMs research requires large data and SAR images are expensive and require specific imaging equipment and algorithms. Privacy and security also prevent the data from becoming open source. Therefore, we are particularly grateful to the publishers of open-source SAR target datasets. By making SARATR-X publicly available, we aim to accelerate the FMs in SAR target recognition by enabling researchers to use our code and weight to design better methods and explore downstream applications.
Figure 8: Visualization of the fourteen datasets included in our pre-training. In this way, with the contribution of open-source researchers and the community, a pre-trained dataset for SAR ATR can be built with multiple targets, scenes, and sensors.
Although this work performs systematical investigations, several limitations and challenges will require exploration in a future study. The SAR images were derived from open-source SAR datasets, and the targets primarily included vehicles, ships, aircraft, oil tanks, etc. Thus, collecting target samples from increasingly unlabeled imagery could further expand the amount of data to reach a million level and the range of downstream applications. The pre-training set image size [123] and automatic target slicing will become important issues that need to be investigated due to the larger image size and contextual range of remote sensing images compared to natural images. In terms of model architecture, learning the dynamic contextual information on space-time is also an important issue. In addition to target shape features, various scenarios and target signatures for SAR self-supervised learning need to be explored. Finally, investigating expert knowledge with text used for multimodal interactions and describing the relationship between targets and scenes could enhance the representation capabilities of FMs for SAR ATR.
In conclusion, we have demonstrated the ability of SARATR-X to adapt to diverse SAR target datasets, achieving high performance and generalizability in classification and detection tasks. By taking full advantage of the rapid growth of SAR images, this SSL-based foundation model opens the door to generalized SAR target recognition.
Appendix A Implementation Details for Section 4 and SARATR-X Experiments